OpenStudy (anonymous):

(a^3b^12c^2) x (a^5c^2) / (b^5c^4)^0 I have no idea how to do this ...

OpenStudy (anonymous):

anything to the power of 0 is 1, so the denominator is just 1

OpenStudy (anonymous):

that leaves \[(a^3b^{12}c^2) \times (a^5c^2) \] which you finish by adding the exponents for the like terms for example \(a^3\times a^5=a^{3+5}=a^8\)

OpenStudy (anonymous):

Ohhh okay. Wow, thank you so much! Definitely helped.

OpenStudy (anonymous):

yw